OpenStudy (anonymous):

negative 1 over 2 x + 1 = –x + 8

OpenStudy (michaelbp):

Simplifying x + 7x = 56 + 8 Combine like terms: x + 7x = 8x 8x = 56 + 8 Combine like terms: 56 + 8 = 64 8x = 64 Solving 8x = 64 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Divide each side by '8'. x = 8 Simplifying x = 8

OpenStudy (michaelbp):

thats just to show you how to do it